OpenStudy (anonymous):

find an equation of the line that passes through (0,7) and is parallel to y=1/4x+4, what is the equation of the line

OpenStudy (anonymous):

The first thing you need to know to solve this problem is that y = 7 x = 0 and m (your slope) = 1/4 in this equation....

OpenStudy (anonymous):

Now, the next thing you have to do, is you have to plug-in those variables... So... 7 = 1/4(0) + b note: b is your y intercept...

OpenStudy (anonymous):

Now you would simplify this to get the value of b... 7 = 0 + b 7 = b Now you can find your answer: y = 1/4x + 7 is parallel to y = 1/4x + 4

OpenStudy (anonymous):

To find the parallel line you could use this equation: y-y1=m(x-x1) So you would plug in your point that is given in the y1 and the x1. So it would be y-7=1/4(x-0) and the final thing you have to do is distribute. your final answer is y=1/4x+7. Remember parallel lines have the same slope :)